OpenStudy (anonymous):

why is this true: 1/2x = x/2 ?

OpenStudy (anonymous):

Cross multiply. 1/2x = x/2  2x(x) = 1 * 2 2x(x) = 2 2(1)(1) = 2 2 = 2 2 times x, times x, is 2, so x has to be 1.